OpenStudy (anonymous):

Find values of a and k for a parabola containing the points given: y - k = a(x + 3)2; (-5,1) and (1,7) A. a = 2; k = 3 B. a = 0.5; k = -1 C. a = .5; k = 3 D. a = 2; k = -1

OpenStudy (anonymous):

\[y=a(x+3)^2+k\] \[1=a(-5+3)^2+k\] \[1=4a+k\] is one equation

OpenStudy (anonymous):

\[7=a(1+3)^2+k\] \[7=16a+k\] is the second equation

OpenStudy (anonymous):

solve \[4a+k=1\]\[16a+k=7\] for \(a\) and \(k\)

OpenStudy (anonymous):

Thanks!